OpenStudy (anonymous):

Can someone help me with this? Will give medal and follow. Question in the comment.

OpenStudy (anonymous):

...?

OpenStudy (anonymous):

\[\frac{ x + 3 }{ 3} = \frac{ 10 + 4 }{ 4 }\]

OpenStudy (anonymous):

Would I put 4(x+3) =3(10+4) and go from there?

OpenStudy (anonymous):

(x+3)*4=14*3 4x+12=42 4x=30 x=30/4

OpenStudy (anonymous):

did you understand

OpenStudy (anonymous):

Yes thank you.